Suppose we have the following annual sales data for an automobile dealership: Year                Sales             &n

  1. Suppose we have the following annual sales data for an automobile dealership:

Year                Sales                Trend

2009                121                     1

2010                187                     2

2011                165                     3

2012                134                     4

2013                155                     5

2014                167                     6

2015                200                     7

2016                206                     8

2017                221                    9

2018                231                 10

We want to forecast sales for 2019 and 2020 using either a simple trend model or a quadratic trend model. Use a within sample forecasting technique to determine the best model using the RMSE measure discussed in lecture. Once this model has been determined, provide actual forecasts for 2019 and 2020.
Answer #1

Since RMSE of quadratic trend model is less than simple linear trend model so quadratic model is better and using this model the forecast for year 2019 and 2020 are given as

using estimated quadratic model, y^=146.18-1.17*t+1.01*t^2

for year 2019, t=11, and y^=146.18-1.17*11+1.01*11*11=255.52

for year 2020, t=12, and y^=146.18-1.17*12+1.01*12*12=277.58

the simple linear trend the model is given as y^=123.93+9.96*t

RMSE=19.42

regression analysis

Observation y y^ y-y^=Residuals(e) e2
1 121 133.89 -12.89 166.18
2 187 143.85 43.15 1862.05
3 165 153.81 11.19 125.30
4 134 163.76 -29.76 885.87
5 155 173.72 -18.72 350.48
6 167 183.68 -16.68 278.18
7 200 193.64 6.36 40.50
8 206 203.59 2.41 5.79
9 221 213.55 7.45 55.48
10 231 223.51 7.49 56.11
sum= 1556 1563.490909 -7.490909091 3769.838
RMSE=sqrt(e2/n)= 19.42
ANOVA
df SS MS F Significance F
Regression 1 8180.148 8180.148 17.10455 0.003272
Residual 8 3825.952 478.2439
Total 9 12006.1
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 123.9333 14.93923 8.295833 3.36E-05 89.48341 158.3833
t 9.957576 2.407674 4.135765 0.003272 4.405469 15.50968

using estimated quadratic model, y^=146.18-1.17*t+1.01*t^2

RMSE=18.07

Observation y y^ y-y^=Residuals(e) e2
1 121 146.03 -25.03 626.36
2 187 147.89 39.11 1529.28
3 165 151.78 13.22 174.68
4 134 157.70 -23.70 561.47
5 155 165.63 -10.63 113.00
6 167 175.59 -8.59 73.75
7 200 187.57 12.43 154.55
8 206 201.57 4.43 19.61
9 221 217.60 3.40 11.58
10 231 235.65 -4.65 21.58
sum= 1556 1551.35 4.65 3264.303
RMSE=sqrt(e2/n)= 18.07
ANOVA
df SS MS F Significance F
Regression 2 8720.217 4360.108 9.288449 0.010724
Residual 7 3285.883 469.4119
Total 9 12006.1
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 146.1833 25.48241 5.736637 0.000708 85.92701 206.4397
t -1.16742 10.64253 -0.10969 0.91573 -26.333 23.99816
t^2 1.011364 0.942888 1.072623 0.31902 -1.21821 3.240939